Can you helf me find the parabola of the form Y=ax^2+bx+c
(0,0),(-4,-36),(2,-18)

Answer by scott8148(6628)

substituting (0,0) __ (0)=a(0)^2+b(0)+c __ 0=c

substituting (-4,-36) __ (-36)=a(-4)^2+b(-4) __ -36=16a-4b

substituting (2,-18) __ (-18)=a(2)^2+b(2) __ -18=4a+2b __ multiplying by 2 __ -36=8a+4b

adding equations __ -36+(-36)=16a-4b+8a+4b __ -72=24a __ dividing by 24 __ -3=a

substituting __ -18=4(-3)+2b __ adding 12 __ -6=2b __ dividing by 2 __ -3=b

Y=-3x^2-3x